three metal cubes with edhes 6 cm 8 cm 10 cm respectively are melt together and formed single cube find volume

Grade:8

1 Answers

Arun
6 years ago
Volume remains same so 
volume of new cube = v1+v2+v3
= 6^3+8^3+10^3
=216+512+1000
=1728
now volume of new cube = a^3= 1728